OpenStudy (anonymous):

Find a and b such that the function f(x)=x2+4x+6 if x<-1 ax+b if x>-1

OpenStudy (anonymous):

\[\begin{cases} & \text{ if } x<-1, x^2 + 4x+6 \\ & \text{ if } x>1, ax+b \end{cases}\] It is probably meant that f(x) has to be both continuous and differentiable (smooth) \[\text{ if } x=-1, x^2 + 4x+6=ax+b\] and \[\text{ if } x=-1, \frac{d(x^2 + 4x+6)}{dx} =\frac{d(ax+b)}{dx}\] Now we evaluate this so \(1 - 4+6 = b-a\), so, \(b-a=3\) and \[\text{ if } x=-1, \frac{d(x^2 + 4x+6)}{dx} =\frac{d( ax+b)}{dx}\] So \(\text{ if } x=-1, 2x + 4 = a\), therefore, \( -2 + 4 = a\) or \(a = 2\) this can be substituted into \(b-a = 3\), so \(b -2= 2\), or \(b = 4\) so we have \(a = 2\), and \(b = 4\)

OpenStudy (anonymous):

@oswaldo12 Sorry I made a typo \(b-a = 3\), hence, \(b-2 = 3\), or \(b = 5\) So we have (a = 2\) and (b = 5\)